At what angle shall a curve be raised in order for a 1500.0-kg vehicle to make a turn around a 100.0-m circular track at a velocity of 30.0 m/s? 




1
Expert's answer
2021-11-30T17:44:58-0500

Solution.

"m=1500.0 kg;"

"r=100.0 m;"

"v=30 m\/s;"

"\theta - ?;"

"Nsin\theta=\dfrac{mv^2}{r};"

"N=\dfrac{mg}{cos\theta};"

"mg\dfrac{sin\theta}{cos \theta}=\dfrac{mv^2}{r};"

"mgtan\theta=\dfrac{mv^2}{r};"

"tan \theta=\dfrac{v^2}{rg};"

"tan \theta=\dfrac{900}{100.0\sdot9.81}=0.9174;"

"\theta=42^o;"

Answer: "\theta=42^o."